User name Sparveris
Log entry time 00:06:52 on May22,2007
Entry number 202906
16:00 Shift takes over with no beam (controlled access)
17:00 Beam permit - trying to get acceptable beam
19:00 Got beam on Be0 @ 5uA (centered on both BPMs).
Run 3689 to check beam profile - looks fine. Changing to 3He target.
19:20 Starting production run 3690 (3He @ 40uA)
19:35 Going to 50uA (during run 3690) to see if we can take it.
19:50 Going to 60uA (during run 3690).
20:00 Going for 70uA (during run 3690).
20:10 MCC had to take us back to 60uA to keep the beam stable.
Starting run 3691 @ 60uA.
20:40 MCC just called - took us up to 70uA (during run 3671) to see if we can hold it.
21:30 Beam down to 30uA (can not give us more)
22:46 Switched to Dummy target - increased raster size by 15%
in both directions. Plan is to take 1 hour of dummy data.
Starting run 3693.
23:28 Dummy run 3694.
00:00 Next shift takes over during run 3694.